Find the length of the diagonal of a rectangle whose sides are 16cm and 12cm
Answer by mananth(16949)

The diagonal and the two sides of rectangle form a right triangle.
The diagonal is the hypotenuse H
16^2 +12^2 = H^2.
H^2 = 400
take square root
H = 20.
The diagonal = 20 cm.